OpenStudy (anonymous):

(AP Physics) Can you please tell me how to figure this out? How do we find the full time of flight of a projectile that travels a distance of 52.5 meters and reaches a maximum height of 30 meters. How can I figure that out?

OpenStudy (anonymous):

didn't we just do this?

OpenStudy (anonymous):

No. Im looking for time. I mean, My way of looking for time is d= Vit + 1/2 gt^2. 30 m = 1/2(9.8)(t^2 So to solve for that I go, 30 / 4.9 = T^2 I need to know if that is correct or not